OBSERVATION OF COHERENT WAVE PACKETS IN THE H^{+}H^{−} ION-PAIR POTENTIAL.

E. REINHOLD, W. HOGERVORST, \underline{W. UBACHS}, Laser Centre, Department of Physics and Astronomy, Vrije Universiteit, Amsterdam, Netherlands.

In an experiment combining 1XUV + 1UV double resonance laser excitation in the presence of a small electric field with pulsed field ionization, bound states in the 1/r potential just below the H^{+}H^{−} ionization (or dissociation) limit were observed. About 10 cm^{−1}below the limit the level density is so high that even for narrowband excitation (Δν_{laser} = 300MHz) about 150 n-states are coherently excited. Experimentally, signal recurrences were observed as a function of delay of the field ionization pulse, that match the splittings between Stark states. A theoretical interpretation of the phenomena is presented fully based on the analysis of Rydberg states in the hydrogen atom, in which the electron is replaced by the heavier H^{−} particle.